OpenStudy (anonymous):

Jaquan buys 3 dozen light bulbs after changing the lightbulbs in his house he has 15 light bulbs left how many lightbulbs did he use?

OpenStudy (anonymous):

21?

OpenStudy (anonymous):

is that a joke if not can you explain you answer

jigglypuff314 (jigglypuff314):

1 dozen = 12 so 3 dozen = 3*12 = 36 then total = (used) + (left) used = (total) - (left) = 36 - 15 = ?

OpenStudy (anonymous):

I love how you help but don't give the answer thats awesome

jigglypuff314 (jigglypuff314):

*tips hat* that's what I do :)